Tuesday, August 22, 2017 - 6:00pm
Client-visible changes from this downtime include:
- Fixed some mistaken user quotas on home directories.
- Upgraded to a 40Gb connection to OARnet.
- Migrated several project directories to new filesets; file 'atime' attributes were not preserved.
- PCON0005, PAS0438, PAS0984, PCON0003, PAS1005, PAS1061, PAS1068, PAS1178, PAS1192, PAS1241, PAS0998, PES0716
- Owens batch server upgrades; potential fix for "server too busy" errors